OpenStudy (anonymous):

Claire goes to a show sale. Everything on the shelf is 35% off the marked price. She falls in love with a pair of red heels that are marked $64.99. If the sales tax rate is 7.5%, what will the total cost be?

OpenStudy (texaschic101):

35% of 64.99 = 0.35(64.99) = 22.746 rounds to 22.75 64.99 - 22.75 = 42.24 --- this is the sale price 42.24 (.075) = 3.168 rounds to 3.17 -- this is the tax 42.24 + 3.17 = 45.41 --- total cost including tax

OpenStudy (anonymous):

thank you :-)))